OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2)

OpenStudy (anonymous):

@kevinbournethird help?

OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2)

OpenStudy (anonymous):

y-8=-0.5(x+17)

OpenStudy (anonymous):

(y + 8) = –2(x –17) (y – 8) =(x + 17) (y – 8) = –2(x +17) (y + 8) =(x – 17)

OpenStudy (anonymous):

thanks :)